OpenStudy (anonymous):

solve x^2 -20=0 with the quadratic formula.... didnt even know this was possible... help!

OpenStudy (campbell_st):

ok... the general form of a quadratic is \[ax^2 + bx + c= 0\] the general quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 1, b = 0 and c = -20 substitute them and evaluate for an answer... you make have to leave it in radical form.